factorize
a^2+b^2+c^2+2ab+2bc+2ca
1
Expert's answer
2012-06-19T08:53:07-0400
a^2+b^2+c^2+2ab+2bc+2ca=(a^2+ab+ac)+(b^2+ab+bc)+(c^2+ac+bc)=a(a+b+c)+b(a+b+c)+c(a+b+c)=(a+b+c)(a+b+c)